David’s basis in the Jimsoo Partnership is $54,500. In a proportionate liquidating distribution, David receives cash of $7,300 and two capital assets: (1) Land A with a fair market value of $20,600 and a basis to Jimsoo of $16,450, and (2) Land B with a fair market value of $10,225 and a basis to Jimsoo of $16,450. Jimsoo has no liabilities.

c1. If the two parcels of land had been inventory to Jimsoo, what are the tax consequences to David (amount and character of gain or loss)?

c2. What is David's basis in distributed assets?

What is David's basis in distributed assets?

Basis
Cash
Land 1
Land 2
Answer #1

SOLUTION:

David's outside basis = 54,500 - (7300 + 16,450 + 16,450) = 14,300

Computation of David’s basis in the distributed assets: David would be allocating the required land increase with unrealized appreciation to an extent of the appreciation. David thus increases the basis of the first parcel land by $4,150 to $20,600. It leaves $10,150 remaining to be allocated

20,600 - 16,450 = 4150

14,300 -4150 = 10,150

David increases it's basis of the two land parcels in proportion of their relative fair market values

Land 1: Basis allocation = 10,150 * 20,600 / (20,600 + 10,225) = 6783

Land 2: Basis allocation = 10,150 * 10,225 / (20,600 + 10,225) = 3367

After completing the allocation David’s bases in the distributed assets are:

Cash: 7300

Land 1: 16,450 + 4150 + 6783 = 27,383

Land 2: 16,450 + 3,367 = 19,817